OpenStudy (anonymous):

A ceramics workshop makes wreaths, trees, and sleighs for sale at Christmas. A wreath takes 3 hours to prepare, 2 hours to paint, and 9 hours to fire. A tree takes 14 hours to prepare, 3 hours to paint, and 4 hours to fire. A sleigh takes 4 hours to prepare, 15 hours to paint, and 7 hours to fire. If the workshop has 116 hours for prep time, 64 hours for painting, and 110 hours for firing, How many of each can be made? A. 8 wreaths, 6 trees, 2 sleighs B. 6 wreaths, 2 trees, 8 sleighs C. 9 wreaths, 7 trees, 3 sleighs D. 2 wreaths, 8 trees, 6 sleighs

HanAkoSolo (jamierox4ev3r):

just try adding the combos. until one adds up to the specified time of 116 hours for prep time, 64 hours for painting, and 110 hours for firing. It would be helpful to use the process of elimination with this

HanAkoSolo (jamierox4ev3r):

@SPrezzy does that make sense?

OpenStudy (anonymous):

I'm still a little lost. Do you mean, for example, for the wreaths use 9+2+3?

HanAkoSolo (jamierox4ev3r):

for each, combine the stats. I will do the first one for you

HanAkoSolo (jamierox4ev3r):

1 Wreath=3 hours prep. 2 hours paint 9 hours fire 1 Tree=14 hours prep. 3 hours paint 4 hours fire 1 Sleigh=4 hours prep. 15 hours paint 7 hours fire

HanAkoSolo (jamierox4ev3r):

I blieve this is correct if i am not mistaken

HanAkoSolo (jamierox4ev3r):

and you have exactly 116 hours prep., 64 hours paint, and 100 fire

HanAkoSolo (jamierox4ev3r):

8(3)+6(14)+2(4)=116 hours prep time 8(2)+6(3)+2(15)=64 hours paint 8(9)+6(4)+2(7)=110 hours fire

HanAkoSolo (jamierox4ev3r):

according to these guidelines @SPrezzy , the first choice happens to be correct! Do you understand how I came to this answer?

OpenStudy (anonymous):

I do. Thank you so much!

HanAkoSolo (jamierox4ev3r):

No problem :)